OpenStudy (anonymous):

Marquette has $150 in his bank account at the beginning of week 1 and will add $25 at the beginning of each subsequent week for tutoring a neighbor in math. What is the explicit formula for this sequence? Write in terms of A(n) and use no spaces in your formula.

OpenStudy (lochana):

Did you try it?

OpenStudy (anonymous):

yes

OpenStudy (lochana):

So what happened :D , alright so n is the number of weeks. and the formula must be in terms of n

OpenStudy (lochana):

When n is 0 which is beginning of the 1 st week(that is week 0) , amount is 150 A(0) = 150 But what about week 1 A(1) = 1*25 + 150 and week 2 A(2) = 2*25 + 150 and so on the formula is A(n) = 25*n + 150

OpenStudy (anonymous):

okay